### v2

# Averaging of Conformity Scores for Classification

Abstract
Conformal prediction provides a robust framework for generating prediction sets with finite-sample coverage guarantees, independent of the underlying data distribution. However, existing methods typically rely on a single conformity score function, which can limit the efficiency and informativeness of the prediction sets. In this paper, we present a novel approach that enhances conformal prediction for multi-class classification by optimally averaging multiple conformity score functions. Our method involves assigning weights to different score functions and employing various data splitting strategies. Additionally, our approach bridges concepts from conformal prediction and model averaging, offering a more flexible and efficient tool for uncertainty quantification in classification tasks. We provide a comprehensive theoretical analysis grounded in Vapnik–Chervonenkis (VC) theory, establishing finite-sample coverage guarantees and demonstrating the efficiency of our method. Empirical evaluations on benchmark datasets show that our weighted averaging approach consistently outperforms single-score methods by producing smaller prediction sets without sacrificing coverage.
1 Introduction
Conformal prediction Vovk et al. (2005); Manokhin (2022) is a robust framework that generates prediction sets with finite-sample coverage guarantees, irrespective of the underlying data distribution Angelopoulos et al. (2023). The fundamental principle of conformal prediction is to construct a prediction set for a new test instance based on the training data, ensuring that the true label is included with a probability of at least $1-\alpha$. This coverage assurance holds regardless of the specific point prediction algorithm employed, making conformal prediction a versatile tool for uncertainty quantification in machine learning. In the split conformal prediction framework Papadopoulos et al. (2002); Lei and Wasserman (2014); Vovk et al. (2018), the training data is partitioned into a training set and a calibration set. The predictive model is trained on the training subset, while the score functions are evaluated on the calibration subset. The conformal prediction set then comprises all labels whose scores fall below a specific quantile of the calibration scores, with the quantile determined by the coverage level.
The choice of score function is critical in determining the efficiency of the resulting prediction sets, especially for multi-class classification. Well-chosen score functions can lead to more informative and precise predictions. This flexibility allows conformal prediction to adapt to the specific characteristics of the data and distribution. Consequently, developing score functions that optimize informativeness and efficiency for various problem settings, including regression Papadopoulos et al. (2008, 2011); Romano et al. (2019); Kivaranovic et al. (2020); Guan (2023); Colombo (2023, 2024) and multi-class classification Sadinle et al. (2019); Romano et al. (2020); Angelopoulos et al. (2021); Huang et al. (2024); Luo and Zhou (2024), remains an active area of research. This work focuses on enhancing the efficiency of prediction sets for classification tasks. While the underlying idea can extend to regression, this paper will focus on the tasks of multi-class classification.
Our approach assumes the availability of multiple score functions for the same classification task, each differing due to variations in the classification algorithm or the definition of the score. We propose assigning optimal weights to aggregate these score functions. Using a validation set, we determine a threshold to achieve the desired coverage and identify the weight combination that minimizes the prediction set size. The final prediction is then based on this weighted score function. Our aim is to find the optimal weights for linear combinations of score functions, thereby fully leveraging the strengths of existing score functions. While our approach shares similarities with Yang and Kuchibhotla (2024), it stands out in three key aspects:
  1. 1.
    Weighted Averaging of Score Functions: Instead of selecting the single best-performing score function, our approach combines multiple score functions through optimal weighting. This averaging can yield more efficient prediction sets than any individual score function while maintaining the desired coverage guarantees.
  2. 2.
    Novel Data Splitting Strategies: We explore and categorize several data splitting methods to determine the optimal weights for combining score functions. In addition to Validity First Conformal Prediction (VFCP) and Efficiency First Conformal Prediction (EFCP) discussed in Yang and Kuchibhotla (2024), we introduce Data Leakage Conformal Prediction (DLCP) and its variant DLCP+, which utilize all available data to enhance weight determination.
  3. 3.
    Theoretical Foundations Using VC Theory: We provide a theoretical analysis of our method that leverages Vapnik–Chervonenkis theory to establish coverage guarantees and expected prediction set sizes. This solid mathematical foundation underscores the validity and efficiency of our approach across different data splitting strategies.
Beyond its novel contributions to the conformal prediction literature, our method is closely related to model averaging Claeskens and Hjort (2008), a well-established technique in machine learning. Unlike traditional model averaging, which assigns weights to different models to improve prediction accuracy, our method assigns weights to score functions. This distinction requires the development of specific data splitting techniques to ensure the desired coverage guarantees. Consequently, our work can be viewed as an innovative adaptation of model averaging principles to the conformal prediction framework.
The remainder of the paper is organized as follows. In Section 3, we detail our weighted averaging approach and the various data splitting strategies employed. Section 3 presents the theoretical analysis, establishing coverage guarantees and expected prediction set sizes. In Section 4, we demonstrate the effectiveness of our method through experiments. Related works are discussed in Section 5. We conclude in Section 6 and outline future research directions.
2 Methodology
2.1 Conformal Prediction for Classification
We start by assuming that a $K$-class classification algorithm provides $\widehat{p}_{y}(x)$, which approximates $P(Y=y|X=x)$ for $y\in[K]$. While our method and theoretical analysis do not depend on the accuracy of this approximation, it is beneficial to assume that higher values of $\widehat{p}_{y}(x)$ indicate a greater likelihood of sample $x$ having label $y$. We consider this training procedure to be performed on a separate dataset, ensuring that $\widehat{p}_{y}(x)$ is independent of the dataset used in this paper.
Let us first consider a single non-conformity score function $s(x,y)$. This function is defined such that higher values of $s(x,y)$ indicate a higher priority of believing $x$ has label $y$. A common choice for the non-conformity score is $s(x,y)=\widehat{p}_{y}(x)$ Sadinle et al. (2019). In this context, conformal prediction for classification problems can be described by the following algorithm. To summarize, we first find a threshold in a set of labeled data, so that $s(x_{i},y_{i})\geq Q_{1-\alpha}$ holds for at least $1-\alpha$ proportion in the set $\mathcal{I}_{2}$. Then we use this threshold to define the prediction set for $x_{i}$ in the test set, which is the upper level set of the function $y\mapsto s(x_{i},y)$.
This algorithm constructs prediction sets $\widehat{C}_{1-\alpha}(x_{i})$ for each $i\in\mathcal{I}_{\textrm{test}}$, based on the non-conformity scores and a threshold determined by the desired coverage probability $1-\alpha$. The algorithm splits the training data into two subsets: $\mathcal{I}_{1}$ for calculating the non-conformity scores, and $\mathcal{I}_{2}$ for calibration. The threshold $Q_{1-\alpha}$ is chosen to ensure the desired coverage probability under the exchangeability assumption.
Algorithm 1 Split Conformal Prediction
1:Data $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$, $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$, pre-determined coverage probability $1-\alpha$
2:A prediction set $\widehat{C}_{\alpha}(x_{i}),i\in\mathcal{I}_{\textrm{test}}$
3:Randomly split $\mathcal{I}_{\textrm{train}}$ into $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$.
4:Train a model $\widehat{p}(x)$ on $\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{1}}$.
5:$Q_{1-\alpha}\leftarrow\lceil(1+|\mathcal{I}_{2}|)(1-\alpha)\rceil$-th smallest non-conformity score $s(x_{i},y_{i})$ for $i\in\mathcal{I}_{2}$.
6:$\widehat{C}_{1-\alpha}(x_{i},Q_{1-\alpha})\leftarrow\{y\in[K]:s(x_{i},y)\geq Q_{1-\alpha}\}$ for $i\in\mathcal{I}_{\textrm{test}}$.
$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\text{train}}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{3}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{1}}$$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{2}}$ $Q_{1-\alpha}^{(1)}(\bm{w})$ by Eq. (2) $\big{(}\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}$ by Eq. (3) $\widehat{\bm{w}}$ by Eq. (4) $Q_{1-\alpha}^{(2)}$ by Eq. (5) $\{x_{i}\}_{i\in\mathcal{I}_{\text{test}}}$ $\left(\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\right)_{i\in\mathcal{I}_{\text{test}}}$ by Eq. (6)
Figure 1: This example illustrates a framework for data splitting into $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$, and $\mathcal{I}_{\text{test}}$. Algorithm 2 presents the complete procedure. Briefly, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are used in Steps 1-2 to select the optimal weight $\widehat{\mathbf{w}}$, while $\mathcal{I}_{3}$ is used in Step 3 as the calibration set for $\mathcal{I}_{\text{test}}$ predictions. We present four options: VFCP, EFCP, DLCP, and DLCP+. Their coverage and size properties are discussed theoretically in Section 3 and empirically in Section 4.
2.2 Various Score Functions for Classifications
In conformal prediction, the choice of score function $s(x,y)$ critically influences the prediction sets. These functions act as non-conformity measures that quantify how appropriately a label matches an input.
Score functions are determined by two primary factors: (1) The predictive model’s quality in estimating $\widehat{p}_{y}(x)\approx\mathbb{P}(Y=y|X=x)$, and (2) The specific methodology for converting these probability estimates into scores. We will consider approaches:
  1. 1.
    Threshold (THR) Sadinle et al. (2019): $s_{\text{THR}}(x,y)=\widehat{p}_{y}(x)$
  2. 2.
    Adaptive Prediction Sets (APS) Romano et al. (2020): $s_{\text{APS}}(x,y)=\sum_{y^{\prime}\in[K]}\widehat{p}_{y^{\prime}}(x)\mathbf{1}\{\widehat{p}_{y^{\prime}}(x)\leq\widehat{p}_{y}(x)\}$
  3. 3.
    Rank-based (RANK) Luo and Zhou (2024): $s_{\text{RANK}}(x,y)=|\{y^{\prime}\in[K]:\widehat{p}_{y^{\prime}}(x)<\widehat{p}_{y}(x)\}|$
These score functions integrate directly with Algorithm 1. While the original formulations include additional refinements to achieve exact $1-\alpha$ coverage guarantees, these refinements do not materially affect our combination methodology and are omitted for simplicity. Our theoretical analysis remains invariant to specific score function design choices. Other notable approaches like RAPS Angelopoulos et al. (2021) and SAPS Huang et al. (2024) demonstrate alternative methodologies for deriving scores from $\widehat{p}_{y}(x)$.
2.3 Averaging Score Functions
In Yang and Kuchibhotla (2024), the authors propose their method with multiple score functions. Their approach selects the score function that yields the smallest average prediction set size among the available score functions. They also introduce two data splitting methods: Efficiency First Conformal Prediction (EFCP) and Validity First Conformal Prediction (VFCP).
Given a vector of score functions $s(x,y)=(s_{1}(x,y),\dots,s_{d}(x,y))^{\top}$, instead of simply choosing the score that provides the smallest prediction set on a calibration set, we propose assigning weights $\bm{w}\in\mathcal{W}\subseteq\mathbb{R}^{d}$ to the non-confomity scores and defining the weighted score function:
$\langle\bm{w},s(x,y)\rangle=\sum_{j=1}^{d}\bm{w}_{j}s_{j}(x,y).$ (1)
By leveraging different scores to create a completely new score function, we expect the weighted score to outperform any individual score. This approach allows for more flexibility in combining the strengths of various score functions, potentially leading to improved prediction set efficiency.
Algorithm 2 Optimal Weight Selection and Prediction Set Construction
1:$\{(x_{i},y_{i})\}_{i\in\mathcal{I}_{\textrm{train}}}$: Labeled training data; $\{x_{i}\}_{i\in\mathcal{I}_{\textrm{test}}}$: Unlabeled test data; $\alpha\in(0,1)$: Pre-defined significance level; $s_{i}(\cdot,\cdot),i=1,\dots,d$: score functions. $\mathcal{W}\subseteq\mathbb{R}^{d}$: Set of possible weight vectors.
2:$\big{(}C_{1-\alpha}(X_{i})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$: Sequence of prediction sets for test data with anticipated $1-\alpha$ coverage.
3:Decide $\mathcal{I}_{1}\subseteq\mathcal{I}_{\textrm{train}}$, $\mathcal{I}_{2}\subseteq\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$, and $\mathcal{I}_{3}\subseteq\mathcal{I}_{\textrm{train}}$ for Step 1, 2, and 3, respectively
4:Step 1: Grid Search over Weights
5:for $\bm{w}\in\mathcal{W}$ do
6:     $Q_{1-\alpha}^{(1)}(\bm{w})\leftarrow$ Calibration($\bm{w},\mathcal{I}_{1},\alpha$).
7:end for
8:Step 2: Optimal Weight Selection
9:for $\bm{w}\in\mathcal{W}$ do
10:     $\big{(}\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))\big{)}_{i\in\mathcal{I}_{2}}$
11:             $\leftarrow$ Evaluation($\bm{w},\mathcal{I}_{2},Q_{1-\alpha}^{(1)}(\bm{w})$).
12:     Compute average prediction set size $S{(\bm{w})}=\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i};\bm{w},Q_{1-\alpha}^{(1)}(\bm{w})\big{)}|$.
13:end for
14:$\widehat{\bm{w}}\leftarrow\arg\min_{\bm{w}\in\mathcal{W}}S{(\bm{w})}$.
15:Step 3: Test-time Prediction Set
16:$Q_{1-\alpha}^{(2)}\leftarrow$ Calibration($\widehat{\bm{w}},\mathcal{I}_{3},\alpha$).
17:$\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$leftarrow Evaluation($\widehat{\bm{w}},\mathcal{I}_{\textrm{test}},\mathcal{I}_{3}$).
18:return $\big{(}\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})\big{)}_{i\in\mathcal{I}_{\textrm{test}}}$
2.4 The Optimal Weight and the Threshold
In this section, we present a detailed procedure for determining the optimal weight vector $\bm{w}$ in the context of the weighted score approach introduced in Eq. (1). This procedure extends Algorithm 1 by incorporating an additional step to minimize the expected size of the prediction set. The following steps outline the process:
  1. 1.
    Extract $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ from dataset. Assume $\mathcal{I}_{1}=\mathcal{I}_{2}$ and $\mathcal{I}_{1},\mathcal{I}_{3}$ partition $\mathcal{I}_{\textrm{train}}$ for clarity.
  2. 2.
    Define $\mathcal{W}\subseteq\mathbb{R}^{d}$ for $\bm{w}$ candidates. For non-negative weights, use grid of simplex $\Delta(d)\subseteq\mathbb{R}^{d}$.
  3. 3.
    Estimate the threshold for every $\bm{w}\in\mathcal{W}$ on $\mathcal{I}_{1}$. In the set $\mathcal{I}_{1}$, calculate the threshold for every $\bm{w}\in\mathcal{W}$:
    $Q_{1-\alpha}^{(1)}(\bm{w})=\ \lceil(1+|\mathcal{I}_{1}|)(1-\alpha)\rceil\text{-th smallest}\text{value of}\quad\{\langle\bm{w},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{1}\}.$ (2)
    This corresponds to Step 5 in Algorithm 1.
  4. 4.
    Determine the prediction set for data in $\mathcal{I}_{2}$. For a given $\bm{w}$ and threshold $Q_{1-\alpha}^{(1)}(\bm{w})$, define the prediction set for each sample $x_{i},i\in\mathcal{I}_{2}$ as:
    $\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))=\{y\in[K]:\langle\bm{w},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(1)}(\bm{w})\}.$ (3)
    This corresponds to Step 6 in Algorithm 1. However, we are not done yet, since these are the prediction sets for all $\bm{w}\in\mathcal{W}$, and our goal is to minimize the prediction set size. This intuitively leads to the next step.
  5. 5.
    The optimal weight vector $\widehat{\bm{w}}$ is obtained by minimizing the empirical prediction set size:
    $\widehat{\bm{w}}\in\operatorname*{arg\,min}_{\bm{w}\in\mathcal{W}}\sum_{i\in\mathcal{I}_{2}}|\widehat{C}_{1-\alpha}(x_{i},\bm{w},Q_{1-\alpha}^{(1)}(\bm{w}))|,$ (4)
    where $\mathcal{W}$ is a discrete grid over the $(d-1)$-dimensional probability simplex $\Delta^{d-1}$. We systematically explore candidate weights via grid search with step size $\varepsilon=0.01$, as detailed in Appendix A.
  6. 6.
    If $\mathcal{I}_{3}\neq\mathcal{I}_{1}$, then $\mathcal{I}_{3}$ is treated as the calibration set in Step 5 in Algorithm 1, and we obtain:
[…]
    value of $\{\langle\widehat{\bm{w}},s(x_{i},y_{i})\rangle:i\in\mathcal{I}_{3}\}.$
    If $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$ are disjoint, $\widehat{\bm{w}}$ and $Q_{1-\alpha}^{(2)}$ are independent. If $\mathcal{I}_{1}=\mathcal{I}_{3}$, then this quantile has been computed in Eq. (2), i.e., $Q_{1-\alpha}^{(2)}=Q_{1-\alpha}^{(1)}(\widehat{\bm{w}})$.
  7. 7.
    The final output is the confidence set for samples $i\in\mathcal{I}_{\textrm{test}}$ in the test set:
    $\widehat{C}_{1-\alpha}(x_{i},\widehat{\bm{w}},Q_{1-\alpha}^{(2)})=\{y\in[K]:\langle\widehat{\bm{w}},s(x_{i},y)\rangle\geq Q_{1-\alpha}^{(2)}\}.$ (6)
This procedure comprises two primary steps: Threshold Calibration: Determining appropriate thresholds using Equations Eq. (2) and Eq. (5) based on the calibration subsets $\mathcal{I}_{1}$ and $\mathcal{I}_{3}$. Prediction Set Evaluation: Constructing the prediction sets for $\mathcal{I}_{2}$ and $\mathcal{I}_{\text{test}}$ as outlined in Equations Eq. (3) and Eq. (6). These steps are encapsulated in Function 1 and Function 2, respectively. The entire procedure of the proposed method is summarized in Algorithm 2.
Function 1 Calibration
1:function Calibration($\bm{w},\mathcal{I}\subseteq\mathcal{I}_{\textrm{train}},\alpha$)
2:     for $i\in\mathcal{I}_{\textrm{calib}}$ do
3:         Compute $\langle\bm{w},s(x_{i},y_{i})\rangle$ using $\bm{w}$ and $(x_{i},y_{i})$.
4:     end for
5:     $Q_{1-\alpha}\leftarrow$ $\lceil(1+|\mathcal{I}|)(1-\alpha)\rceil$-th smallest score $\langle\bm{w},s(x_{i},y_{i})\rangle$ using $\bm{w}$ and $(x_{i},y_{i})$.
6:     return $Q_{1-\alpha}$.
7:end function
Function 2 Evaluation
1:function Evaluation($\bm{w},\mathcal{I}\subseteq\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}},Q_{1-\alpha}$)
2:     for $i\in\mathcal{I}$ do
3:         for $y\in[K]$ do
4:              Compute weighted score $\langle\bm{w},s(x_{i},y)\rangle$
5:         end for
6:         $\widehat{C}_{1-\alpha}(x_{i})\leftarrow\{y\in[K]:\langle\bm{w},s(x_{i},y)\rangle\geq Q_{1-\alpha}\}$
7:     end for
8:     return $\big{(}\widehat{C}_{1-\alpha}(x_{i})\big{)}_{i\in\mathcal{I}}$
9:end function
2.5 Data Splitting
In Algorithm 2, the method for splitting the data into $\mathcal{I}_{1}$, $\mathcal{I}_{2}$, and $\mathcal{I}_{3}$ has not been specified. To explore potential data splitting approaches, we first highlight the following two key observations. Firstly, After determining $\widehat{\bm{w}}$, the calibration procedure Eq. (5) and prediction set construction Eq. (6) mirror Step 5 and Step 6 of Algorithm 1. Prior steps aim to identify a weight vector $\bm{w}$ optimizing algorithm performance. Secondly, Eq. (2) and Eq. (5) find quantiles, requiring sample labels. Equations Eq. (3) and Eq. (6) define prediction sets, needing only feature $x$. Thus, the test set can be included in Eq. (3) and Eq. (6).
Based on the observations above, we introduce the following four possible ways of data splitting.
  • (a)
    Validity First Conformal Prediction (VFCP) Yang and Kuchibhotla (2024): $\mathcal{I}_{1}=\mathcal{I}_{2}\subseteq\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}\setminus\mathcal{I}_{1}$. Divides $\mathcal{I}_{\textrm{train}}$ into two partitions. Marginal coverage probability is guaranteed to be at least $1-\alpha$ under exchangeability of samples in $\mathcal{I}_{3}$ and $\mathcal{I}_{\textrm{test}}$.
  • (b)
    Efficiency First Conformal Prediction (EFCP) Yang and Kuchibhotla (2024): $\mathcal{I}_{1}=\mathcal{I}_{2}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$. Uses all training data to determine $\widehat{\bm{w}}$, resulting in more accurate estimation of optimal $\bm{w}$. This method equires stronger assumptions for coverage guarantee.
  • (c)
    Data Leakage Conformal Prediction (DLCP): $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$, and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$. Minimizes prediction set on $\mathcal{I}_{\textrm{test}}$ in finding $\widehat{\bm{w}}$ in Eq. (4). Called ”data leakage” as it uses test data in training procedure of $\widehat{\bm{w}}$.
  • (d)
    Data Leakage Conformal Prediction+ (DLCP+): $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$. Uses all available data in each step, including $\mathcal{I}_{\textrm{test}}$ in $\mathcal{I}_{2}$ to maximize sample size for finding $\widehat{\bm{w}}$ in Eq. (4).
There are various possible methods for data splitting. In the following sections, we focus on these four specific approaches, examining their theoretical properties and evaluating their performance through experiments.
3 Theoretical Analysis
3.1 Overview of the Results
We investigate the theoretical properties of the proposed methods in terms of validity and efficiency.
Validity: We assess whether the output prediction set achieves the desired coverage rate of $1-\alpha$. For the VFCP method, the coverage rate is guaranteed under the exchangeability assumption. However, for the other three methods, establishing validity is more complex due to the selection bias introduced by $\widehat{\bm{w}}$. Consequently, our aim is to demonstrate that the optimization over $\bm{w}\in\mathcal{W}$ has small impact on validity.
Efficiency: We first observe that the prediction set attains the smallest possible expected size only if the true conditional probabilities $p(y|x)$ are known. Given the predefined score functions, the optimal weight vector in the population sense is defined as
$\bm{w}^{*}=$ $\ \arg\min_{\bm{w}\in\mathcal{W}}\min_{q\in\mathbb{R}}\mathbb{E}\left[|\widehat{C}_{1-\alpha}(X,\bm{w},q)|\right]$
s.t. $\quad\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q)\right)\geq 1-\alpha.$
$\bm{w}^{*}$ is the weight vector that minimizes the expected prediction set size while ensuring that, together with an appropriate threshold $q$, the prediction set maintains a desired coverage rate of $1-\alpha$. Equivalently, we can define
$Q_{1-\alpha}(\bm{w}):=\sup\{q\in\mathbb{R}:\mathbb{P}(Y\in\widehat{C}_{1-\alpha}(X,\bm{w},q))\geq 1-\alpha\},$
and using this definition, we define
$\bm{w}^{*}\in\arg\min_{\bm{w}\in\mathcal{W}}\mathbb{E}\left[|\widehat{C}_{1-\alpha}(X,\bm{w},Q_{1-\alpha}(\bm{w}))|\right].$
We define $q^{*}=Q_{1-\alpha}(\bm{w}^{*})$. Since $\bm{w}^{*}$ and $q^{*}$ are deterministic once $\alpha$ and the true distribution are given, we can succinctly denote the optimal prediction set for $x$:
$\widehat{C}_{1-\alpha}^{*}(x):=\widehat{C}_{1-\alpha}(x,\bm{w}^{*},q^{*}).$
Our analysis of efficiency focuses on the difference between the size of the prediction set produced by the proposed methods using $\widehat{\bm{w}}$ and the expected size of $\widehat{C}^{*}_{1-\alpha}(x)$. In other words, we analyze the discrepancy between the prediction set sizes generated by $\widehat{\bm{w}}$ and the optimal $\bm{w}^{*}$. We will demonstrate that this difference is negligible given a sufficiently large dataset.
We note that this analysis differs from verifying whether $\widehat{\bm{w}}$ converges to $\bm{w}^{*}$, which would require assumptions about the smoothness of the function mapping the weight to the prediction set size. Instead, our analysis makes fewer assumptions on the optimal prediction set size.
3.2 Results by Vapnik–Chervonenkis Theory
We begin by defining the following events on the probability space of pairs of exchangeable random variables $(X_{i},Y_{i})_{i\in\mathcal{I}}$. The event $\Omega(\mathcal{I},\eta)$ is defined as:
$\begin{split}\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}&\Big{|}\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathbf{1}\{\langle\bm{w},s(X_{i},Y_{i})\rangle\geq q\}\\ &-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,Y)\rangle\geq q\}]\Big{|}\leq\eta,\end{split}$ (7)
where $(X,Y)$ has the same distribution as $(X_{1},Y_{1})$. $\Gamma(\mathcal{I},\xi)$ denotes the event
$\begin{split}\sup_{\bm{w}\in\mathbb{R}^{d},q\in\mathbb{R}}\Big{|}&\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(X_{i},y)\rangle\geq q\}\\ &-\mathbb{E}[\mathbf{1}\{\langle\bm{w},s(X,y)\rangle\geq q\}]\Big{|}\leq\xi,\end{split}$ (8)
where we recall that $K$ is the number of classes of the classification task. $\Omega(\mathcal{I},\eta)$ is about the uniform concentration of the coverage, and $\Gamma(\mathcal{I},\xi)$ is about the uniform concentration of the prediction set size. This is because by the definition of $\widehat{C}_{1-\alpha}$ in Eq. (2) or Eq. (5),
[…]
$\sum_{y\in[K]}\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq q\}=\big{|}\widehat{C}_{1-\alpha}(x,\bm{w},q)\big{|}.$
The following lemma shows that these two events hold with high probability.
Lemma 1.
Suppose the samples in $\mathcal{I}$ are i.i.d., then
  • (a)
    $\Omega\left(\mathcal{I},8\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+\delta\right)$ hold with probability at least $1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right)$.
  • (b)
    $\Gamma\left(\mathcal{I},8K\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+K\delta\right)$ hold with probability at least $1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right)$.
The proof of this lemma appears in Section B in the appendix. This lemma establishes that $\Omega(\mathcal{I},\eta)$ and $\Gamma(\mathcal{I},\xi)$ hold with high probability provided that $\eta,\xi\gtrsim\sqrt{\frac{d\log|\mathcal{I}|}{|\mathcal{I}|}}$. The proof, presented in the appendix, applies subgraph classes from Vapnik-Chervonenkis (VC) theory. Additionally, if the inequalities in Equations Eq. (7) and Eq. (8) hold for $\bm{w}\in\mathbb{R}^{d}$, they naturally extend to $\bm{w}\in\mathcal{W}\subset\mathbb{R}^{d}$.
Remark.
Part (b) of Lemma 1 limits the generalization of our current theoretical results to classification tasks. Compared to the bound in Equation Eq. (7), the bound in Equation Eq. (8) involves a summation over $[K]$. In the proof, this is handled using a union bound. However, for regression tasks where prediction set size is typically measured by the Lebesgue measure, a union bound over $K$ is not directly applicable. Therefore, our current work focuses exclusively on classification tasks. Extending the proposed method to regression tasks is left for future research.
3.3 Consistency of VFCP
The statistical guarantees of VFCP requires a few additional assumptions, e.g., the continuity between the quantile and the prediction set. To get rid off such assumptions and present the result in a neater way, we will only prove a modified version of VFCP. We will change $\alpha$ in Eq. (2) to a slightly smaller $\alpha^{\prime}$. The condition for $\alpha^{\prime}$ will be specified in the theorem.
Theorem 1.
Let $\widehat{C}_{1-\alpha}^{\text{VFCP}}(x)$ be the output of Algorithm 2 with the setting of VFCP, and modified as mentioned above. Suppose the samples in $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$ are exchangeable, then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\right)\geq 1-\alpha.$
Moreover, if $\Omega(\mathcal{I}_{1},\eta_{1}),\Omega(\mathcal{I}_{3},\eta_{3})$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ are satisfied, and $\alpha^{\prime}+\eta_{1}+\eta_{3}\leq\alpha$, then for test sample $X$,
$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{VFCP}}(X)\big{|}\right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{2},$
where $1-\alpha_{1}=\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha^{\prime})\rceil-\eta_{3}.$
The proofs of all theorems appear in Section C and Section D and the appendix.
The validity of VFCP requires minimal assumption (exchangeability), while it is less efficient than other method. We will verify this fact in empirical studies.
3.4 Consistency of EFCP
In the setting of EFCP, $\mathcal{I}_{1}=\mathcal{I}_{2}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$. We denote this set by $\mathcal{I}_{\textrm{train}}$.
Theorem 2.
Let $\widehat{C}_{1-\alpha}^{\text{EFCP}}(x)$ be the output of Algorithm 2 with the setting of EFCP. Suppose the data in $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ are independent and $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ is satisfied. Then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}^{\text{EFCP}}(X)\right)\geq 1-\alpha_{1}\quad\text{where}$
11 11 $-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
Moreover, if $\Gamma(\mathcal{I}_{\textrm{train}},\xi_{\text{train}})$ is satisfied, then for any $X$ in the test set,
$\mathbb{E}\left[\big{|}\widehat{C}_{1-\alpha}^{\text{EFCP}}(X)\big{|}\right]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+2\xi_{\text{train}}.$
3.5 Consistency of DLCP
In the setting of DLCP, $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$. We will denote this by $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ respectively. In this case, the test set and $\widehat{\bm{w}}$ become dependent. The prediction set size is presented differently than previous results.
Theorem 3.
Let $\widehat{C}_{1-\alpha}^{\text{DLCP}}(x))$ be the output of Algorithm 2 for a test sample $x$ with the setting of DLCP. Suppose $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion satisfies
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}$ $\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP}}(x_{i})\}\geq 1-\alpha_{1}-\eta_{\text{test}}\quad\text{where}$
$1-\alpha_{1}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
In addition, if $\Gamma(\mathcal{I}_{\textrm{test}},\xi_{\text{test}})$ holds, then
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}|\widehat{C}_{1-\alpha}^{\text{DLCP}}(x_{i})|\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha_{1}}(X)\big{|}\right]+\xi_{\text{test}}.$
3.6 Consistency of DLCP+
Similar as the DLCP setting, $\widehat{\bm{w}}$ depends on the test set, so the result of prediction set size is presented in a similar way. In the following theorem, we use $\mathcal{I}_{2}$ to denote $\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$.
Theorem 4.
Let $\widehat{C}_{1-\alpha}^{\text{DLCP+}}(x,\widehat{\bm{w}}))$ be the output of Algorithm 2 with the setting of DLCP+. Suppose $\Omega(\mathcal{I}_{\textrm{train}},\eta_{\text{train}})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}_{1-\alpha}^{\text{DLCP+}}(x_{i},\widehat{\bm{w}})\}\geq 1-\alpha^{\prime}-\eta_{\text{test}}$
$\text{where}\quad 1-\alpha^{\prime}=\frac{1}{|\mathcal{I}_{\textrm{train}}|}\lceil(1+|\mathcal{I}_{\textrm{train}}|)(1-\alpha)\rceil-\eta_{\text{train}}.$
In addition, if $\Gamma(\mathcal{I}_{2},\xi_{2})$ holds, then for test sample $X$,
$\mathbb{E}[|\widehat{C}_{1-\alpha}^{\text{DLCP+}}(X)|]\leq\mathbb{E}\left[\big{|}\widehat{C}^{*}_{1-\alpha^{\prime}}(X)\big{|}\right]+2\xi_{2}.$
3.7 Conclusion of Theoretical Results
We integrate the results of Lemma 1 with our main theorems. The Lemma 1 indicates that as the size of the set $\mathcal{I}$ approaches infinity, the events $\Omega(\mathcal{I},\eta)$ and $\Gamma(\mathcal{I},\xi)$ hold with probability $1-o(1)$ even if $\eta,\xi\to 0$ is at a sufficient slow rate. Consequently, the terms $\eta_{i}$ and $\xi_{i}$ in our theorems become negligible. This suggests that the proposed methods achieve coverage rates close to $1-\alpha$ and exhibit near-optimal efficiency.
4 Experiments
4.1 Score Weighting
We conducted an experiment to compare the performance of various score functions score functions discussed in Section 2.2: THR Sadinle et al. (2019), APS Romano et al. (2020), and RANK Luo and Zhou (2024) given by a pretrained classifier $\widehat{p}_{y}(x)$. Additional details about implementation can be found in Section E in the appendix. Throughout the experiment, we assumed that a pretrained classifier $\widehat{p}_{y}(x)$ was available, and the split of the dataset unseen during the training of the pretrained classifier.
Figure 2: Boxplot comparison of different score functions at a significance level of $\alpha=0.01$ on CIFAR-100. Our weighted combination method achieves the guaranteed coverage of 99% while maintaining the smallest prediction set size.
In the experiments on CIFAR-10 and CIFAR-100, testing images, which were not used during the pretraining of the model, were used as the $\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{\textrm{test}}$ sets. The experiments were performed for different significance levels $\alpha$ ranging from 0.01 to 0.05. 100 runs with different index splits were conducted to ensure robustness. We have additional experiments on data splitting ratio, which can be found in Section F in the supplimentary file.
Figure 3: Comparison of size vs. coverage for various score functions and our proposed method across $\alpha$ values (0.01-0.05). Our weighted combination method (red) consistently outperforms the other baseline methods by achieving the desired coverage rate with smaller prediction set sizes.
$\alpha=0.01$ $\alpha=0.05$
Method Coverage Size Coverage Size
VFCP 0.990 (0.003) 13.782 (1.114) 0.950 (0.005) 3.890 (0.266)
EFCP 0.989 (0.003) 13.306 (0.464) 0.949 (0.005) 3.754 (0.096)
DLCP 0.989 (0.003) 13.298 (0.461) 0.949 (0.005) 3.752 (0.097)
DLCP+ 0.989 (0.003) 13.299 (0.459) 0.949 (0.005) 3.753 (0.097)
APS 0.990 (0.003) 40.217 (1.786) 0.949 (0.006) 19.545 (1.022)
THR 0.989 (0.003) 26.949 (2.198) 0.949 (0.006) 4.519 (0.381)
RANK 0.990 (0.003) 27.161 (2.421) 0.950 (0.006) 6.298 (0.395)
RAPS 0.990 (0.003) 26.403 (2.505) 0.950 (0.005) 5.581 (0.263)
SAPS 0.990 (0.003) 25.096 (2.298) 0.950 (0.006) 5.036 (0.302)
Table 1: Coverage and size of different methods for $\alpha=0.01$ and $\alpha=0.05$ on CIFAR-100 dataset. Results are shown as mean (standard deviation). The first four methods corresponds to the splitting methods in Section 2.5.
The primary objective of our experiments was to evaluate the performance of the proposed weighted score function in comparison with three foundational base score functions: APS, THR, and RANK. Additionally, we compared our method against two competitive baseline score functions, RAPS Angelopoulos et al. (2021) and SAPS Huang et al. (2024).
Figure 2 compares coverage and prediction set sizes across methods at $\alpha=0.01$ on CIFAR-100, while Figure 3 evaluates performance across $\alpha\in[0.01,0.05]$. Using VFCP splits for consistency, our method achieves guaranteed coverage with the smallest prediction sets overall. The advantages are particularly pronounced on CIFAR-10 across all $\alpha$ values and on CIFAR-100 for $\alpha\leq 0.02$.
Notably, THR exhibits strong performance on less challenging tasks (CIFAR-10), where our weighting scheme naturally assigns it dominant weights. In these cases, score averaging provides limited improvement. However, for complex multi-class scenarios (CIFAR-100) or stringent significance levels ($\alpha<0.05$), our method demonstrates clear superiority by optimally combining score functions.
Furthermore, to explore the influence of different data splitting strategies (as discussed in Section 2.4), we conducted additional comparisons of our weighted score function using various split approaches. We also included the performance results of the five individual score functions when utilizing the VFCP split method. Table 1 provides a comprehensive summary of the coverage and size metrics at significance levels of $\alpha=0.01$ and $\alpha=0.05$. These results highlight the effectiveness of the different data splitting strategies. It is important to note that while APS aims to achieve conditional coverage, the other methods, including our approach, do not specifically target conditional coverage. Consequently, it is not surprising that the alternative methods exhibit better efficiency compared to APS.
4.2 Model Weighting
In addition to choosing the score functions in Section 2.2, we further conducted an experiment to compare the performance of various models and their weighted combinations for prediction set construction on CIFAR-10 and CIFAR-100. For CIFAR-10, we used the models ResNet-56, ShuffleNetV2 (1.0x), and VGG16-BN. For CIFAR-100, we used the models VGG16-BN, RepVGG-A2, and MobileNetV2 (1.0x). The performance of the weighted combination method was compared against individual models.
Specifically, for each of the four score functions: THR, APS, RAPS, and SAPS, we generated prediction sets by combining the scores of different models using weights selected by our proposed method. The results are visualized in Figures 4 and 5, which demonstrate the effectiveness of our weighted combination method in terms of weighing scores of different models.
Figure 4: Across various score functions, our weighted combination of models outperformed any individual model and achieved optimal size on the CIFAR-10 dataset across $\alpha$ values (0.01–0.05).
Figure 5: Across various score functions, our weighted combination of models outperformed any individual model and achieved optimal size on the CIFAR-100 dataset across $\alpha$ values (0.01–0.05).
5 Related Work
Our work builds upon existing research in conformal prediction, particularly focusing on model averaging and calibration techniques. In the realm of model aggregation, Yang et al. Yang and Kuchibhotla (2024) introduced two selection algorithms designed to minimize the width of prediction intervals by aggregating and selecting from multiple regression estimators.
The technique of VC dimension has been applied in Yang and Kuchibhotla (2024) and Candès et al. (2023). In Section D.2 of Yang and Kuchibhotla (2024), the uniform probability bound is over a finite set. In Candès et al. (2023), the authors considers maximization over real numbers. They have not consider the techiniques subgraph theory like our approach. The methods proposed in Carlsson et al. (2014); Linusson et al. (2017) can aggregate different prediction models to perform conformal prediction, but the methods do not focus on optimizing the efficiency.
A concurrent preprint Liang et al. (2024) examines parametrized score functions $s_{\lambda}(x,y)$ for $\lambda\in\Lambda$ in regression settings. While their general framework handles arbitrary parameter spaces, controlling the Rademacher complexity becomes challenging except in specialized cases. Our approach differs fundamentally by combining pre-defined score functions linearly. This structural constraint enables theoretical guarantees dependent only on the number of constituent scores rather than the complexity of function spaces.
6 Conclusion and Discussion
Our weighted score aggregation method enables efficient and valid prediction set construction for multi-class classification through optimized combinations of score functions and strategic data splitting. Theoretically, we establish finite-sample coverage guarantees and oracle inequalities quantifying the efficiency gap between our method and optimal weights. Empirically, experiments demonstrate consistent maintenance of coverage requirements with minimal prediction set sizes compared to single-score baselines. This work bridges model averaging and conformal prediction, providing a flexible framework for uncertainty quantification that adapts to dataset characteristics through optimal score combinations.
These results suggest several directions for future work:
  1. 1.
    Regression Extension: Algorithmic adaptation to regression requires new theoretical tools due to continuous output spaces, where finite-class union bounds become inapplicable. Potential approaches include metric entropy analysis or covering number techniques.
  2. 2.
    Optimization Enhancement: Developing gradient-based alternatives to grid search, such as differentiable conformal objectives or online weight adaptation during model training, would improve scalability with many score functions.
Appendix A Grid Search Implementation
To solve Eq. (4), we discretize the probability simplex:
$\Delta^{d-1}=\left\{\bm{w}\in\mathbb{R}_{+}^{d}:\sum_{j=1}^{d}w_{j}=1\right\}$
using a grid resolution $\varepsilon=0.01$. Candidate weights are generated as:
$\mathcal{W}=\left\{\bm{w}=(k_{1}\varepsilon,\ldots,k_{d}\varepsilon)\,\,\bigg{|}\,\,k_{j}\in\mathbb{N},\,\,\sum_{j=1}^{d}k_{j}=\lfloor 1/\varepsilon\rfloor\right\},$
yielding $\binom{\lfloor 1/\varepsilon\rfloor+d-1}{d-1}$ distinct weight vectors. For $d=3$ scores, this produces 5,151 candidates, ensuring comprehensive coverage of the parameter space while remaining computationally tractable through parallel evaluation.
Appendix B Proof of Lemma 1
B.1 VC Dimension of the Subgraph Classes
Proposition 1.
Both of the following classes of functions
$\{(x,y)\mapsto\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq t\}:\bm{w}\in\mathbb{R}^{d},t\in\mathbb{R}\}$ (9)
and
$\{x\mapsto\mathbf{1}\{\langle\bm{w},s(x,y)\rangle\geq t\}:\bm{w}\in\mathbb{R}^{d},t\in\mathbb{R}\},$ (10)
where $y\in[K]$ is fixed, have VC-dimension at most $d+1$.
Proof.
Both of the classes are the subgraph classes of vector space of functions with dimension $d+1$. By Proposition 4.20 of Wainwright (2019), these subgraph classes have dimension at most $d+1$. This is also a direct result of Example 4.21 in the book. ∎
B.2 Proof of the Lemma
This proof basically follows from Vapnik–Chervonenkis theory. We will use the theorems and lemmas in Wainwright (2019) as reference. The proof of both parts of the lemma are almost identical. We will focus on part (a). Let $\epsilon_{i},i\in\mathcal{I}$ be i.i.d. symmetric random variables takes value $-1$ or $1$, i.e., $\mathbb{P}(\epsilon_{i}=-1)=\mathbb{P}(\epsilon_{i}=1)=0.5.$ Then we define the Rademacher complexity for the function class in Eq. (9),
$\mathcal{R}(\mathcal{I}):=\mathbb{E}\left[\sup_{w,t}\left|\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\epsilon_{i}\mathbf{1}\{\langle w,s(X_{i},Y_{i})\rangle\geq t\}\right|\right].$
By Theorem 4.10 in the book,
$\mathbb{P}\left(\Omega\left(\mathcal{I},\mathcal{R}(\mathcal{I})+\delta\right)\right)\geq 1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right).$ (11)
By Proposition 1, the VC dimension of the function class in Eq. (9) is $d+1$. Then altogether with Lemma 4.14 and Proposition 4.18 in the book,
$\mathcal{R}(\mathcal{I})\leq 4\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}$ (12)
We combine Eq. (11) and Eq. (12) to obtain the result of part (a). The proof of part (b) of the lemma follows analogously, using the VC dimension result for the function class in Eq. (10). Let us define
$\Gamma_{y}(\mathcal{I},\eta)=\sup_{w\in\mathbb{R}^{d},t\in\mathbb{R}}\left|\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\mathbf{1}\{\langle w,s(X_{i},y)\rangle\geq t\}-\mathbb{E}[\mathbf{1}\{\langle w,s(X,y)\rangle\geq t\}]\right|\leq\eta.$
For $\mathcal{I}\subseteq\mathcal{I}_{\textrm{train}}\cup\mathcal{I}_{\textrm{test}}$,
$\mathbb{P}\left(\Gamma_{y}\left(\mathcal{I},8\sqrt{\frac{(d+1)\log(|\mathcal{I}|+1)}{|\mathcal{I}|}}+\delta\right)\right)\geq 1-\exp\left(-\frac{|\mathcal{I}|\delta^{2}}{2}\right).$
It is clear that
$\bigcup_{y\in[K]}\Gamma_{y}(\mathcal{I},\eta)\subseteq\Gamma(\mathcal{I},K\eta).$
Taking the union bound on $y\in[K]$ implies the result of part (b).
Appendix C Preliminary Lemmas for the Theorems
C.1 Bounds for Coverage Probability
The following result requires the data for optimizing $\widehat{\bm{w}}$ are independent with the calibration set and the test set. The algorithm satisfying these conditions have the most reliable coverage probability.
Lemma 2.
Let $\widehat{C}_{1-\alpha}(x)$ be the output of Algorithm 2. Suppose
  • (i)
    $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$ are exchangeable.
  • (ii)
    $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{1}\cup\mathcal{I}_{2}}$ and $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$ are independent.
Then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}(Y\in\widehat{C}_{1-\alpha}(X))\geq 1-\alpha.$
Proof.
$\widehat{\bm{w}}$ only depends on $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{1}\cup\mathcal{I}_{2}}$, so it is independent of $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$. The weighted score function $\{\langle\widehat{\bm{w}},s(X_{i},Y_{i})\rangle\}_{i\in\mathcal{I}_{3}\cup\mathcal{I}_{\textrm{test}}}$ are also exchangeable. For any $(X,Y)$ in the test set, the rank of $s(X,Y)$ is smaller than $Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$ with probability $\frac{\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil}{1+|\mathcal{I}_{3}|}\geq 1-\alpha$. ∎
Lemma 3.
Let $\widehat{C}_{1-\alpha}(x)$ be the output of Algorithm 2. Suppose
  • (i)
    $\Omega(\mathcal{I}_{3},\eta_{3})$ is satisfied.
  • (ii)
    $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{1}\cup\mathcal{I}_{2}}$ and $\{(X_{i},Y_{i})\}_{i\in\mathcal{I}_{\textrm{test}}}$ are independent.
Then for $(X,Y)$ in the test set, the coverage probability
$\mathbb{P}(Y\in\widehat{C}_{1-\alpha}(X))\geq\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil-\eta_{3}.$
Proof.
Let us write the threshold in Eq. (5) $Q_{1-\alpha}^{(2)}:=Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$ to emphasize that this quantile depends on $\widehat{\bm{w}}$ and the samples in $\mathcal{I}_{3}$. By the procedure of the algorithm, we have
$\mathbb{P}\left(Y\in\widehat{C}_{1-\alpha}(X)\right)=\mathbb{P}\left(\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\right).$
Assuming the event $\Omega(\mathcal{I}_{3},\eta_{3})$,
$\begin{split}Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})&=\sup\left\{t\in\mathbb{R}:\frac{1}{|\mathcal{I}_{3}|}\sum_{i\in\mathcal{I}_{3}}\mathbf{1}\{\langle\widehat{\bm{w}},s(x_{i},y_{i})\rangle\geq t\}\geq\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil\right\}\\ &\geq\sup\left\{t\in\mathbb{R}:\mathbb{P}(\langle\widehat{\bm{w}},s(X,Y)\rangle\geq t)\geq\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil-\eta_{3}\right\}\\ &=Q_{1-\alpha^{\prime}}(\widehat{\bm{w}}),\end{split}$ (13)
where $1-\alpha^{\prime}=\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil-\eta_{3}.$ By the definition of the quantile,
$\mathbb{P}(\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3}))\geq\mathbb{P}(\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha^{\prime}}(\widehat{\bm{w}}))=1-\alpha^{\prime}.$ (14)
The proof is complete. ∎
Lemma 4.
Let $\widehat{C}_{1-\alpha}(X;\widehat{\bm{w}})$ be the output of Algorithm 2. Suppose $\Omega(\mathcal{I}_{3},\eta_{3})$ and $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$ hold, then the coverage proportion satisfies
$\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{y_{i}\in\widehat{C}(x_{i})\}\geq\frac{1}{|\mathcal{I}_{3}|}\lceil(1+|\mathcal{I}_{3}|)(1-\alpha)\rceil-\eta_{3}-\eta_{\text{test}}.$
Proof.
Under the event $\Omega(\mathcal{I}_{\textrm{test}},\eta_{\text{test}})$, for all $w\in\mathcal{W}$,
$\left|\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},Y_{i})\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\}-\mathbb{E}_{(X,Y)}[\mathbf{1}\{\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\}]\right|$
$\leq\sup_{\bm{w}\in\mathbb{R}^{d},t\in\mathbb{R}}\left|\frac{1}{|\mathcal{I}_{\textrm{test}}|}\sum_{i\in\mathcal{I}_{\textrm{test}}}\mathbf{1}\{\langle\bm{w},s(X_{i},Y_{i})\rangle\geq t\}-\mathbb{E}_{(X,Y)}[\mathbf{1}\{\langle\bm{w},s(X,Y)\rangle\geq t\}]\right|\leq\eta_{\text{test}},$
where $(X,Y)$ is an i.i.d. copy of $(X_{i},Y_{i})$ in the test set. In particular, we let $\bm{w}=\widehat{\bm{w}}$ and $t=Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$, we have
$E[\mathbf{1}\{\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha^{\prime}}(\widehat{\bm{w}},\mathcal{I}_{3})\}\mid\widehat{\bm{w}}\ ]=\mathbb{P}(\langle\widehat{\bm{w}},s(X,Y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\mid\ \widehat{\bm{w}})=1-\alpha.$
The remaining proof has similar argument as Eq. (13) and Eq. (14), and is omitted here. ∎
C.2 Bounds for Prediction Set Size
Lemma 5.
Suppose the samples in $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ satisfy $\Omega(\mathcal{I}_{1},\eta_{1})$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ respectively, and suppose $Q_{1-\alpha^{\prime}}(\widehat{\bm{w}},\mathcal{I}_{1})\leq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$, then for $X$ in the test set, the expected prediction set size satisfies
$\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\}\leq\mathbb{E}\left[\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha_{1}}(w^{*})\}\right]+\xi_{2},$
where $1-\alpha_{1}=\frac{1}{|\mathcal{I}_{1}|}\lceil(1+|\mathcal{I}_{1}|)(1-\alpha^{\prime})\rceil+\eta_{1}$.
Proof.
Under the assumption $Q_{1-\alpha^{\prime}}(\widehat{\bm{w}},\mathcal{I}_{1})\leq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})$, for $w\in\mathcal{W}$, we have
$\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w,s(X_{i},y)\rangle\geq Q_{1-\alpha}(w,\mathcal{I}_{3})\}\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w,s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w,\mathcal{I}_{1})\}.$
$\widehat{\bm{w}}$ is obtained from
$\widehat{\bm{w}}\in\arg\min_{w\in\mathcal{W}}\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w,s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w,\mathcal{I}_{1})\}$
Therefore,
$\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha^{\prime}}(\widehat{\bm{w}},\mathcal{I}_{1})\}\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w^{*},\mathcal{I}_{1})\}.$
Given the event $\Omega(\mathcal{I}_{1},\eta_{1})$, for $w\in\mathcal{W}$ and $t\in\mathbb{R}$,
$\mathbb{P}(\langle w,s(X,Y)\rangle\geq t)=\mathbb{E}[\mathbf{1}\{\langle w,s(X,Y)\rangle\geq t\}]\leq\frac{1}{|\mathcal{I}_{1}|}\sum_{i\in\mathcal{I}_{1}}\mathbf{1}\{\langle w,s(X,Y)\geq t\}+\eta_{1}.$
The lower bound of the LHS is also a lower bound of the RHS, so we have
$\begin{split}Q_{1-\alpha^{\prime}}(w^{*},\mathcal{I}_{1})&=\sup\left\{t\in\mathbb{R}:\frac{1}{|\mathcal{I}_{1}|}\sum_{i\in\mathcal{I}_{1}}\mathbf{1}\{\langle w,s(x_{i},y_{i})\rangle\geq t\}\geq\frac{1}{|\mathcal{I}_{1}|}\lceil(1+|\mathcal{I}_{1}|)(1-\alpha^{\prime})\rceil\right\}\\ &\geq\sup\left\{t\in\mathbb{R}:\mathbb{P}(\langle w^{*},s(X,Y)\rangle\geq t)\geq\frac{1}{|\mathcal{I}_{1}|}\lceil(1+|\mathcal{I}_{1}|)(1-\alpha^{\prime})\rceil+\eta_{1}\right\}\\ &=Q_{1-\alpha_{1}}(w^{*}),\end{split}$ (15)
where $1-\alpha_{1}=\frac{1}{|\mathcal{I}_{1}|}\lceil(1+|\mathcal{I}_{1}|)(1-\alpha^{\prime})\rceil+\eta_{1}$ and $Q_{1-\alpha_{1}}(\bm{w}^{*})$ is the $(1-\alpha_{1})$-quantile of the population distribution $\langle w,s(X,Y)\rangle$. Now we can conclude that
$\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w^{*},\mathcal{I}_{1})\}\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha_{1}}(w^{*})\}.$
On the event $\Gamma(\mathcal{I}_{2},\xi_{2})$,
$\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w^{*})\}\leq\mathbb{E}\left[\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha_{1}}(w^{*})\}\right]+\xi_{2}.$
The proof is complete. ∎
Appendix D Proof of the Theorems
D.1 Proof of Theorem 1
Under the assumption of the theorem, the setting of VFCP satisfies the condition of Lemma 2. This proves the coverage probability in the theorem. For the expected prediction set size, under the event $\Gamma(\mathcal{I}_{2},\xi_{2})$, by Lemma 5, for $X$ in the test set,
$\mathbb{E}\left[\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\bm{w},\mathcal{I}_{3})\}\right]\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\bm{w},\mathcal{I}_{3})\}+\xi_{2}.$
This implies the coverage probability of the theorem.
D.2 Proof of Theorem 2
Under the assumption of the theorem, the setting of VFCP satisfies the condition of Lemma 3. This proves the coverage probability in the theorem. For the expected prediction set size, since $\mathcal{I}_{1}=\mathcal{I}_{3}$ and $\Gamma(\mathcal{I}_{2},\xi_{2})$ is satisfied, by Lemma 5, for $X$ in the test set,
$\mathbb{E}\left[\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\bm{w},\mathcal{I}_{3})\}\right]\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\bm{w},\mathcal{I}_{3})\}+\xi_{2}.$
This implies the coverage probability of the theorem.
D.3 Proof of Theorem 3
Since $\mathcal{I}_{1}=\mathcal{I}_{3}=\mathcal{I}_{\textrm{train}}$ and $\mathcal{I}_{2}=\mathcal{I}_{\textrm{test}}$, one can verify that the theorem is the direct result of Lemma 4 and Lemma 5.
D.4 Proof of Theorem 4
Similar as the setting of DLCP, the coverage probability is the direct result of Lemma 4. For the prediction set size, suppose $\Gamma(\mathcal{I}_{2},\xi_{2})$ holds, then for $X$ in the test set,
$\mathbb{E}[|\widehat{C}_{1-\alpha}^{\text{DLCP+}}(X)|]$ $\leq\frac{1}{|\mathcal{I}_{2}|}\sum_{i\in\mathcal{I}_{2}}\sum_{y\in[K]}\mathbf{1}\{\langle\widehat{\bm{w}},s(X_{i},y)\rangle\geq Q_{1-\alpha}(\widehat{\bm{w}},\mathcal{I}_{3})\}+\xi_{2}$
$\leq\mathbb{E}\left[\sum_{y\in[K]}\mathbf{1}\{\langle w^{*},s(X,y)\rangle\geq Q_{1-\alpha^{\prime}}(w^{*})\}\right]+2\xi_{2}.$
Appendix E Additional Discussion on Score Functions
The three basic score functions used in our experiment were:
Least Ambiguous Set Values Classifier (THR) Sadinle et al. (2019).
The score function of THR is defined as:
$s_{\text{THR}}(x,y)=\widehat{p}_{y}(x).$
This score function is straightforward: it assigns a higher score to labels with a higher estimated probability. The prediction set includes labels with the highest estimated probabilities. This is the score function we have to include because if $\widehat{p}$ is the true posterior probability, then the score function itself can achieve the smallest expected prediction set size.
Adaptive Prediction Set (APS) Romano et al. (2020).
The score function of APS is defined as:
$s_{\text{APS}}(x,y)=\sum_{y^{\prime}\in[K]}\widehat{p}_{y^{\prime}}(x)\mathbf{1}\{\widehat{p}_{y^{\prime}}(x)\leq\widehat{p}_{y}(x)\},$
where $\mathbf{1}\{\cdot\}$ is the indicator function. This score function can be interpreted as the complement of the $p$-value for label $k$. It measures the sum of the estimated probabilities of all labels that have the same or a smaller estimated probability than label $k$.
Rank-based Score Function (RANK) Luo and Zhou (2024).
The score function of RANK is defined as:
$s_{\text{RANK}}(x,y)=\frac{|\{k^{\prime}\in[K]:\widehat{p}_{k^{\prime}}(x)<\widehat{p}_{k}(x)\}|}{K-1},$
This score function assigns a score based on the rank of the estimated probability $\widehat{p}_{y}(x)$ among all the estimated probabilities for input $x$. The rank is divided by $K-1$ so that the range of the score is from 0 to 1. The prediction set gives higher priority to labels with larger ranks.
The three score functions, THR, APS, and RANK, employ different strategies for assigning scores based on the estimated probabilities. THR directly utilizes the estimated probabilities as scores. APS, on the other hand, considers the cumulative probability of labels that have the same or lower estimated probabilities compared to the label of interest. RANK, in contrast, focuses on the relative ranking of the estimated probabilities among all possible labels. It is important to note that all these score functions preserve the order of the labels, which means that, for fixed $x$, the order of labels based on their estimated probabilities, $\widehat{p}_{y}(x)$, remains the same when ranked according to their scores, $s(x,y)$.
Appendix F Additional Experiments with Different Data Splitting Ratio
(a) Comparison of coverage and size for different data split methods at a significance level of $\alpha=0.05$ when $\mathcal{I}_{\textrm{train}}:\mathcal{I}_{\textrm{test}}$ = 80:20. EFCP, DLCP, and DLCP+ exhibit similar size results, but DLCP has the smallest coverage and the largest gap from the desired coverage level of $0.951-\alpha=$ 0.95. VFCP attains the desired coverage at the cost of having the largest prediction set size.
(b) Comparison of coverage and size for different data split methods at a significance level of $\alpha=0.05$ when $\mathcal{I}_{\textrm{train}}:\mathcal{I}_{\textrm{test}}$ = 99:1. DLCP achieves the smallest size among the compared methods. VFCP attains the desired coverage at the cost of having the largest prediction set size.
Figure 6: Comparison of coverage and size for different data split methods at $\alpha$=0.05.